# Stable self-adaptive timestepping for Reduced Order Models for incompressible flows

### v1

@@ Abstract
This work introduces RedEigCD, the first self-adaptive timestepping technique specifically tailored for reduced-order models (ROMs) of the incompressible Navier-Stokes equations. Building upon linear stability concepts, the method adapts the timestep by directly bounding the stability function of the employed time integration scheme using exact spectral information of matrices related to the reduced operators. Unlike traditional error-based adaptive methods, RedEigCD relies on the eigenbounds of the convective and diffusive ROM operators, whose computation is feasible at reduced scale and fully preserves the online efficiency of the ROM.
A central theoretical contribution of this work is the proof, based on the combined theorems of Bendixson and Rao, that, under linearized assumptions, the maximum stable timestep for projection-based ROMs is shown to be larger than or equal to that of their corresponding full-order models (FOMs).
@@ 1 Introduction
Beyond timestep control at the FOM level, several contributions have addressed adaptive or stability-aware strategies in the reduced-order modeling framework. Amsallem and Farhat [18] proposed interpolation and stabilization techniques for parametric ROMs, emphasizing the importance of maintaining spectral consistency of the reduced operators. Their work highlighted the importance of maintaining consistency between the ROM and FOM eigenvalues to ensure numerical robustness. Nevertheless, this stabilization strategy does not explicitly link timestep adaptivity to linear stability theory, nor does it exploit the exact spectral information that can be computed efficiently at a reduced scale. Rezaian and Wei [19] demonstrated that analyzing and reassigning the eigenvalues of the reduced system’s linear operator provides an effective means to control instability while preserving dominant dynamical features, establishing spectral analysis as a central tool for stability-aware model reduction.
Even though lots of efforts are devoted to pushing the boundaries of POD-Galerkin ROMs applied to CFD (closure models [20, 21], structure-preserving operators [22, 23, 24], hyper-reduction [25, 26, 24, 23], stabilization [27, 28, 29], etc.), little effort has been made to finetune the timestep to maximize the ”performance” of the time integration for these ROMs. Sanderse [8] indicated the use of the Gershgorin circle theorem as a possible way to estimate the eigenbounds of the combined reduced system to set the timestep, however no tests or implementation were presented. For structural mechanics, Bach et al. [30] performed an extensive analysis of the stability conditions for projection-based ROMs, where the maximum stable timesteps for both FOM and ROM were compared analytically. It was observed that the maximum stable timestep at the ROM level, assuming a linear problem, would not be smaller than the maximum stable timestep at the FOM level. To the best of our knowledge, such an analysis has not been performed for the nonlinear incompressible Navier-Stokes equations.
Therefore, this work aims to bridge the gap between self-adaptive timestepping techniques developed for scale-resolving CFD simulations [6, 7] and reduced-order models, as according to Bach et al. [30], it should be possible to generally obtain larger timesteps. Thus, the potential benefit of using a ROM is twofold: the reduced cost due to dimension reduction, and marching faster in time due to larger timesteps.
@@
The present work introduces a new class of stability-driven self-adaptive timestepping algorithms for reduced-order models of the incompressible Navier-Stokes equations. The key novelty lies in (i) extending existing self-adaptive timestep control strategies - initially developed for full-order CFD solvers - to projection-based ROMs while fully preserving reduced-order efficiency, and (ii) establishing a theoretical proof, based on the combined theorems of Bendixson [32] and Rao [33], that the maximum stable timestep of a ROM is always larger or equal to that of its FOM counterpart. This connection between linear stability theory and reduced-order modeling has not been reported before for nonlinear incompressible flow systems. Furthermore, the proposed algorithm, termed RedEigCD, achieves stability control without resorting to full-order model operations. Lastly, it applies to both periodic and inflow-outflow configurations. Together, these elements form a novel stability-aware time-integration framework that advances the state-of-the-art in efficient and reliable ROM-based CFD simulations.
@@ 2.1 Full-order model
where $C$ is the convective operator, $D$ is the diffusive operator, $G$ is the gradient operator, $\mathbf{f}$ is the discrete forcing term, and $\Omega$ is a diagonal matrix with the volumes of the cells as entries; $\mathbf{u},\mathbf{p}$ are the velocity and kinematic pressure. Note that this formulation can be easily adapted to methods that are not the finite-volume method. The exact form of these operators depends on the details of the chosen discretization. Within the framework of model order reduction, Eqs. (3a)-(3b) are identified as the FOM.
@@
where $\tilde{F}(\mathbf{u})=\Omega^{-1}(-C(\mathbf{u})+D)\mathbf{u}+\mathbf{f}$.
From this point onwards, the discretization of the operators is assumed to be performed in a symmetry-preserving manner for the sake of simplicity. Symmetry-preserving schemes [35, 34] are a family of schemes that preserve the symmetry properties of the continuous operators at the discrete level. This is usually achieved by ensuring that the discrete gradient and divergence operators are dual to each other, i.e., $\Omega M^{T}=G$; the convective operator is discretized in a skew-symmetric manner, and the diffusive operator is discretized as a symmetric matrix and corresponds to the composition of the gradient and divergence operators, resulting in a negative semi-definite matrix.
@@ 2.2 Linear stability analysis for the FOM
where $F=\Omega^{-1}(D-C(\mathbf{u}))\in\mathbb{R}^{m\times m}$ is the operator arising from linearizing $\tilde{F}(\mathbf{u})$ in Eq.(4), and $R(z)$ is the stability (or augmentation) function of the given (linearized) scheme. Note that there is an approximation symbol instead of an equal symbol as the $n+1$ value requires projecting every stage in the set of incompressible velocity fields [37, 38], which is left out in Eq.(5). However, the projection step in a projection method is implicit and does not affect the stability region.
@@
The methodology explained in this subsection assumes the discretization is symmetry-preserving. The reader is referred to Appendix B of [7] to generalize the method to non-symmetry-preserving schemes, such as upwinding schemes, by splitting the convective operator in its symmetric and skew-symmetric contributions.
As explained in Section 2.1, the symmetry-preserving discretization ensures that certain property from the continuous operators are preserved [35]: a symmetry-preserving diffusive operator is represented by a real semi-negative definite symmetric matrix, and the convective operator is represented by a real skew-symmetric matrix [35, 34]. Hence, the eigenvalues of the former lie on the negative real axis in the complex plane, while the eigenvalues of the latter lie on the imaginary axis. As eigenbounds are subadditive, i.e. $\rho(A+B)\leq\rho(A)+\rho(B)$, where $\rho(A)$ determines the eigenbounds of the matrix $A$, the eigenbound of the operator $F$, $\lambda_{F}$, lies, according to Bendixson’s inequality [32], in the rectangle defined by the eigenbounds of the convective and diffusive operators as shown in Figure 1.
@@ 2.4 Energy-conserving POD-Galerkin method
The ROM considered in this work follows the ODE-based projection approach [42], where the PDE is first discretized in space, as in Eq. (3), and afterwards is projected to the subspace obtained with POD [8].
Let the velocity vector be defined as $\mathbf{u}\in\mathbb{R}^{N_{V}}$, and the POD-Galerkin approximation is obtained by the Ansatz that the velocity field can be approximated by
@@
where the weighted orthonormality condition has been applied. $\mathbf{a}$ stands for the convected velocity in the reduced space while $\mathbf{a}_{c}=\mathbf{a}$ stands for the convecting velocity in the reduced space. Applying the duality of the divergence and gradient operators inherent in a symmetry-preserving discretization, the pressure term is exactly zero, and can be dropped if the snapshots are divergence-free [8], as shown by
@@
which has a complexity of $\mathcal{O}(M^{3})$. In many cases, the required number of modes $M$ is low enough so that performing operations of this complexity is acceptable. For larger $M$, hyperreduction techniques such as the discrete empirical interpolation method might be used [23, 24, 25] that reduce the complexity to $\mathcal{O}(M^{2})$.
@@
where $\Pi$ interpolates the fluxes to the staggered faces, and $A$ interpolates the velocities to the staggered faces. As described in [44], the boundary contributions $\mathbf{y}_{*},\,*\in\{A,D,\Pi,G,M\}$ can be expressed as linear functions of a vector $\mathbf{y}_{bc}$ that fully describes the prescribed boundary conditions, $\mathbf{y}_{*}=B_{*}\mathbf{y}_{bc}$. The boundary condition vector is approximated as $\mathbf{y}_{bc}(t)\approx\Phi_{bc}\mathbf{a}_{bc}(t)$, where $\mathbf{a}_{bc}\in\mathbb{R}^{M_{bc}}$, and $M_{bc}$ being the number of modes used for the boundary conditions. Inserting this approximation and the Ansatz (15) with $\mathbf{u}_{\mathrm{inhom}}(t)=F_{\mathrm{inhom}}\mathbf{a}_{bc}(t)$ into (17) (projected onto $\Phi$), leads to
@@ 3.1 Estimating the eigenbound of the reduced operators
The concept behind AlgEigCD is approximating the eigenbounds of the operators by finding a factorization $H=AB$ such that the eigenbound of $A^{T}B^{T}$ can be approximated efficiently. The reduced operators, however, have the form $H=\Phi^{T}H_{F}$, where $H_{F}$ is a matrix with FOM dimensions. Following AlgEigCD, the eigenbound of a matrix $H$ would be obtained by approximating the eigenbounds of $H^{*}=(\Phi^{T}A_{F})^{T}(B_{F}\Phi)^{T}$, where $A_{F}B_{F}$ is a factorization of $H_{F}$. In contrast to $H$, $H^{*}$ is a matrix of FOM dimension, therefore the eigenbound estimation for $H^{*}$ is more expensive than for $H$ itself.
Thus, such a transfer does not result in an efficient method. Instead, we will propose a new approach that exploits the structure of the reduced operators (13) and (23).
Concerning the diffusive operator, the application of the exact calculation of the eigenvalues is straightforward, as the diffusive operator at the ROM level is built as $\Phi^{T}D\Phi$, and it is a real symmetric matrix. Thus, its eigenvalues are real, and the eigenbound $\rho(D_{r})$ can be computed exactly in an offline stage.
On the other hand, the convective operator at the ROM level is built as in Eq. (13). Simplifying the momentum equation as a particular case of a convection-diffusion equation with predefined convecting velocity, the Jacobian of the convective operator with respect to the convected velocity coefficients reads as
@@ 3.2 Analytical stability analysis of ROMs
In practice, it has been observed that ROMs are stable for larger timesteps than the FOM. For instance, Baiges et al. [48] indicated that for explicit time integration schemes, the timestep was not restricted by the CFL condition, opposite to the FOM. Intuitively, cutting off some of the highest (spatial) frequencies, which are represented by the smallest singular values, would imply removing some of the most rapid (temporal) modes. These are the modes that make the system stiffer, and thus are the most restrictive modes in terms of stability. By eliminating these modes, the stability constraint is more relaxed, and the method yields, as a consequence, larger timesteps in the reduced framework.
Bach et al. [30] analyzed the stability of ROMs compared to FOMs for linear elasticity problems, which are linear problems with real symmetric matrices (like diffusion), where the timestep at the ROM level is demonstrated to be at least as large as the timestep at the FOM level, yet usually larger.
To provide a mathematical foundation to this qualitative point of view, Bach et al. [30] made use of the Poincaré separation theorem, which reads as follows:
@@
As $A$ is a normal matrix, it can be diagonalized with a unitary matrix $U$, which is a set of orthonormal eigenvectors such that $U^{*}U=I$, and a diagonal matrix $\Lambda$, which has the eigenvalues of $A$ as entries, i.e. $\Lambda=\mathrm{diag}(\lambda_{1},\dots,\lambda_{n})$ [51, pp. 101–103]. Therefore,
@@
However, Theorem 1 serves as a strong theoretical proxy for the stability behavior of the ROM. Since the projection $\Phi$ inherently removes the high-frequency mdoes associated with the largest eigenvalues of $C(\mathbf{u})$, the inequality $\rho(C_{r})<\rho(C)$ is expected to hold robustly in practice. The theorem thus provides the structural justification for the increased stability limits observed in numerical experiments, even in the presence of trajectory divergence.

### v2

@@ Abstract
This work introduces RedEigCD, the first self-adaptive timestepping technique specifically tailored for reduced-order models (ROMs) of the incompressible Navier-Stokes equations that employ stucture-preserving discretizations. Building upon linear stability concepts, the method adapts the timestep by directly bounding the stability function of the employed time integration scheme using exact spectral information of matrices related to the reduced operators. Unlike traditional error-based adaptive methods, RedEigCD relies on the eigenbounds of the convective and diffusive ROM operators, whose computation is feasible at reduced scale and fully preserves the online efficiency of the ROM. A central theoretical contribution of this work is the proof, based on the combined theorems of Bendixson and Rao, that, under linearized assumptions, the maximum stable timestep for projection-based ROMs is shown to be larger than or equal to that of their corresponding full-order models (FOMs).
@@
{highlights}
Introduces RedEigCD, a stability-based self-adaptive timestepper for ROMs.
Provides a theoretical proof that ROMs permit larger stable timesteps than FOMs.
Extends Bendixson and Rao theorems to analyze ROM stability bounds.
Achieves up to 40x larger stable timesteps without accuracy loss.
Establishes a general, stable, efficient framework for adaptive ROM time integration.
@@ 1 Introduction
Beyond timestep control at the FOM level, several contributions have addressed adaptive or stability-aware strategies in the reduced-order modeling framework. Amsallem and Farhat [18] proposed interpolation and stabilization techniques for parametric ROMs, emphasizing the importance of maintaining spectral consistency of the reduced operators. Their work highlighted the importance of maintaining consistency between the ROM and FOM eigenvalues to ensure numerical robustness. Nevertheless, this stabilization strategy does not explicitly link timestep adaptivity to linear stability theory, nor does it exploit the exact spectral information that can be computed efficiently at a reduced scale. Building on the concept of using the eigenvalues of the linear operator to guide subspace rotation [19], Rezaian and Wei [20] demonstrated that analyzing and reassigning the eigenvalues of the reduced system’s linear operator provides an effective means to control instability while preserving dominant dynamical features, establishing spectral analysis as a central tool for stability-aware model reduction.
Even though lots of efforts are devoted to pushing the boundaries of POD-Galerkin ROMs applied to CFD (closure models [21, 22], structure-preserving operators [23, 24, 25], hyper-reduction [26, 27, 25, 24], stabilization [28, 29, 30], etc.), little effort has been made to finetune the timestep to maximize the ”performance” of the time integration for these ROMs. Sanderse [8] indicated the use of the Gershgorin circle theorem as a possible way to estimate the eigenbounds of the combined reduced system to set the timestep, however no tests or implementation were presented. While recent non-intrusive frameworks have employed Gershgorin-based regularization to ensure the stability of learned operators, e.g., Gkimisis et al. [31], the present work realizes Sanderse’s vision [8] in an intrusive context by implementing a stability-based adaptive timestepper. This aligns with efforts in other fields, such as structural mechanics, where Bach et al. [32] performed an extensive analysis of the stability conditions for projection-based ROMs, where the maximum stable timesteps for both FOM and ROM were compared analytically. It was observed that the maximum stable timestep at the ROM level, assuming a linear problem, would not be smaller than the maximum stable timestep at the FOM level. To the best of our knowledge, such an analysis has not been performed for the nonlinear incompressible Navier-Stokes equations.
Therefore, this work aims to bridge the gap between self-adaptive timestepping techniques developed for scale-resolving CFD simulations [6, 7] and reduced-order models, as according to Bach et al. [32], it should be possible to generally obtain larger timesteps. This observation is further supported by recent developments across various ROM paradigms. Beyond machine-learning-augmented frameworks that utilize coarsened timestepping [33, 34, 35], both classical Galerkin [36] and non-intrusive projections [37] have demonstrated the ability to maintain stability and qualitative physics at temporal resolutions significantly coarser than the FOM. Thus, the potential benefit of using a ROM is twofold: the reduced cost due to dimension reduction, and marching faster in time due to larger timesteps.
@@
The present work introduces a new class of stability-driven self-adaptive timestepping algorithms for reduced-order models of the incompressible Navier-Stokes equations. The key novelty lies in (i) extending existing self-adaptive timestep control strategies - initially developed for full-order CFD solvers - to projection-based ROMs while fully preserving reduced-order efficiency, and (ii) establishing a theoretical proof, based on the combined theorems of Bendixson [39] and Rao [40], that the maximum stable timestep of a ROM is always larger or equal to that of its FOM counterpart. This connection between linear stability theory and reduced-order modeling has not been reported before for nonlinear incompressible flow systems. Furthermore, the proposed algorithm, termed RedEigCD, achieves stability control without resorting to full-order model operations. Lastly, it applies to both periodic and inflow-outflow configurations. While the current derivation assumes a structure-preserving, discretely divergence-free discretization, the underlying spectral bounding framework remains generalizable to non-divergence-free formulations through standard operator splitting. Together, these elements form a novel stability-aware time-integration framework that advances the state-of-the-art in efficient and reliable ROM-based CFD simulations.
@@ 2.1 Full-order model
where $C$ is the convective operator, $D$ is the diffusive operator, $G$ is the gradient operator, $M$ is the divergence operator, $\mathbf{f}$ is the discrete forcing term, and $\Omega$ is a diagonal matrix with the volumes of the cells as entries; $\mathbf{u}=\mathbf{u}(t),\mathbf{p}=\mathbf{p}(t)$ are the velocity and kinematic pressure. Note that this formulation can be easily adapted to methods that are not the finite-volume method. The exact form of these operators depends on the details of the chosen discretization. Within the framework of model order reduction, Eqs. (3a)-(3b) are identified as the FOM.
@@
where $\tilde{F}(\mathbf{u},t)=\Omega^{-1}(-C(\mathbf{u})+D)\mathbf{u}+\mathbf{f}$.
From this point onwards, the discretization of the operators is assumed to be performed in a symmetry-preserving manner for the sake of simplicity. Symmetry-preserving schemes [42, 41] are a family of schemes that preserve the symmetry properties of the continuous operators at the discrete level. This is usually achieved by ensuring that the diffusive operator is discretized as a symmetric matrix and corresponds to the composition of the gradient and divergence operators, resulting in a negative semi-definite matrix, the convective operator corresponds to a skew-symmetric matrix, and the discrete gradient and divergence operators are dual to each other, i.e., $\Omega M^{T}=G$. In physical terms, this duality ensures that the pressure-gradient work and the velocity divergence cancel exactly in the global energy balance, preventing artificial energy sources or sinks in the discrete system [42].
@@ 2.2 Linear stability analysis for the FOM
where $F=\Omega^{-1}(D-C(\mathbf{u}))\in\mathbb{R}^{m\times m}$ is the operator arising from linearizing $\tilde{F}(\mathbf{u})$ in Eq.(4) via Picard linearization, and $R(z)$ is the stability (or augmentation) function of the given (linearized) scheme. Note that there is an approximation symbol instead of an equal symbol as the $n+1$ value requires projecting every stage in the set of incompressible velocity fields [44, 45], which is left out in Eq.(5). However, the projection step in a projection method is implicit and does not affect the stability region in a zero-divergence framework.
@@
While the incompressible Navier-Stokes equations are inherently nonlinear, the ultimate goal of this analysis is to ensure numerical stability. This approach is justified by the fact that numerical stability is fundamentally a local phenomenon. By performing a Taylor expansion of the nonlinear spatial operator around the current state $\mathbf{u}^{n}$, $\dot{\mathbf{u}}\approx F(\mathbf{u}^{n})+\frac{\partial F}{\partial\mathbf{u}}(\mathbf{u}_{n})(\mathbf{u}-\mathbf{u}^{n})$, it is evident that the error growth is dictated by the eigenvalues of the linearized operator rather than the global nonlinear system. This justifies that these methods are restricted to linear stability analysis instead of introducing nonlinear methods.
The methodology explained in this subsection assumes the discretization is symmetry-preserving. The reader is referred to Appendix B of [7] to generalize the method to non-symmetry-preserving schemes, such as upwinding schemes, by splitting the convective operator in its symmetric and skew-symmetric contributions. As explained in Section 2.1, the symmetry-preserving discretization ensures that certain property from the continuous operators are preserved [42]: a symmetry-preserving diffusive operator is represented by a real semi-negative definite symmetric matrix, and the convective operator is represented by a real skew-symmetric matrix [42, 41]. Hence, the eigenvalues of the former lie on the negative real axis in the complex plane, while the eigenvalues of the latter lie on the imaginary axis. As eigenbounds are subadditive, i.e. $\rho(A+B)\leq\rho(A)+\rho(B)$, where $\rho(A)$ determines the eigenbounds of the matrix $A$, the eigenbound of the operator $F$, $\lambda_{F}$, lies, according to Bendixson’s inequality [39], in the rectangle defined by the eigenbounds of the convective and diffusive operators as shown in Figure 1.
@@ 2.4 Energy-conserving POD-Galerkin method
The ROM considered in this work follows the ODE-based projection approach [49], where the PDE is first discretized in space, as in Eq. (3), and afterwards is projected to the subspace obtained with POD [8]. Let the velocity vector be defined as $\mathbf{u}\in\mathbb{R}^{N_{V}}$, and the POD-Galerkin approximation is obtained by the Ansatz that the velocity field can be approximated by
@@
where the weighted orthonormality condition has been applied. $\mathbf{a}$ stands for the convected velocity in the reduced space while $\mathbf{a}_{c}=\mathbf{a}$ stands for the convecting velocity in the reduced space. While physically identical in the momentum equation, this notation highlights the functional roles as the state vector and the linearization point, respectively. This separation ensures the framework remains generalizable to non-autonomous transport equations (see C), where the convecting field may be a prescribed velocity. Applying the duality of the divergence and gradient operators inherent in a symmetry-preserving discretization, the pressure term is exactly zero, and can be dropped if the snapshots are divergence-free [8], as shown by
@@
which has a complexity of $\mathcal{O}(M^{3})$ [22]. In many cases, the required number of modes $M$ is low enough so that performing operations of this complexity is acceptable. For larger $M$, hyperreduction techniques such as the discrete empirical interpolation method might be used [24, 25, 26] to reduce the complexity to $\mathcal{O}(M^{2})$. While hyperreduction specifically targets this complexity reduction, for flows at high Reynolds numbers, closure models, e.g., [35], can also be employed. These models ensure stability at very low values of M by modeling the dissipative effects of truncated scales, offering a complementary approach to maintain ROM feasibility.
@@
where $\Pi$ interpolates the fluxes to the staggered faces, and $A$ interpolates the velocities to the staggered faces. Figure 2 represents a staggered grid, where the velocities are calculated at the faces of the collocated grid (gray), and the staggered faces are represented in blue for the $x$-velocity component and in red for the $y$-velocity component.
ppppppppuuuuuuuuuuuuvvvvvvvvvvvv
Figure 2: Representation of a two-dimensional staggered grid, where the collocated grid is depicted in gray, the $x$-staggered grid in blue, and the $y$-staggered grid in red.
As described in [51], the boundary contributions $\mathbf{y}_{*},\,*\in\{A,D,\Pi,G,M\}$ can be expressed as linear functions of a vector $\mathbf{y}_{bc}$ that fully describes the prescribed boundary conditions, $\mathbf{y}_{*}=B_{*}\mathbf{y}_{bc}$. The boundary condition vector is approximated as $\mathbf{y}_{bc}(t)\approx\Phi_{bc}\mathbf{a}_{bc}(t)$, where $\mathbf{a}_{bc}\in\mathbb{R}^{M_{bc}}$, and $M_{bc}$ being the number of modes used for the boundary conditions. Inserting this approximation and the Ansatz (15) with $\mathbf{u}_{\mathrm{inhom}}(t)=F_{\mathrm{inhom}}\mathbf{a}_{bc}(t)$ into (17) (projected onto $\Phi$), leads to
@@ 3.1 Estimating the eigenbound of the reduced operators
The concept behind AlgEigCD is approximating the eigenbounds of the operators by finding a factorization $H=AB$ such that the eigenbound of $A^{T}B^{T}$ can be approximated efficiently. The reduced operators, however, have the form $H=\Phi^{T}H_{F}$, where $H_{F}$ is a matrix with FOM dimensions. Following AlgEigCD, the eigenbound of a matrix $H$ would be obtained by approximating the eigenbounds of $H^{*}=(\Phi^{T}A_{F})^{T}(B_{F}\Phi)^{T}$, where $A_{F}B_{F}$ is a factorization of $H_{F}$. In contrast to $H$, $H^{*}$ is a matrix of FOM dimension, therefore the eigenbound estimation for $H^{*}$ is more expensive than for $H$ itself. Thus, such a transfer does not result in an efficient method. Instead, we will propose a new approach that exploits the structure of the reduced operators (13) and (23).
Concerning the diffusive operator, the application of the exact calculation of the eigenvalues is straightforward, as the diffusive operator at the ROM level is built as $\Phi^{T}D\Phi$, and it is a real symmetric matrix. Thus, its eigenvalues are real, and the eigenbound $\rho(D_{r})$ can be computed exactly in an offline stage. On the other hand, the convective operator at the ROM level is built as in Eq. (13). Simplifying the momentum equation as a particular case of a convection-diffusion equation with predefined convecting velocity, the Jacobian of the convective operator with respect to the convected velocity coefficients reads as
@@
Even though Eq.(26) only holds for skew-symmetric convective operators, it can be generalized to cases that do not fulfill this condition, i.e., upwinding schemes, by splitting each $C_{r,i}$ into its symmetric and skew-symmetric contributions similarly to Eq.(27), and therefore applying the same exact method for the symmetric (real) and skew-symmetric (imaginary) contributions.
@@ 3.2 Analytical stability analysis of ROMs
In practice, it has been observed that ROMs are stable for larger timesteps than the FOM. For instance, Baiges et al. [55] indicated that for explicit time integration schemes, the timestep was not restricted by the CFL condition, as the Courant number could be larger than unity, opposite to the FOM. Intuitively, cutting off some of the highest (spatial) frequencies, which are represented by the smallest singular values, would imply removing some of the most rapid (temporal) modes. These are the modes that make the system stiffer, and thus are the most restrictive modes in terms of stability. By eliminating these modes, the stability constraint is more relaxed, and the method yields, as a consequence, larger timesteps in the reduced framework.
Bach et al. [32] analyzed the stability of ROMs compared to FOMs for linear elasticity problems, which are linear problems with real symmetric matrices (like diffusion), where the timestep at the ROM level is demonstrated to be at least as large as the timestep at the FOM level, yet usually larger. To provide a mathematical foundation to this qualitative point of view, Bach et al. [32] made use of the Poincaré separation theorem, which reads as follows:
@@
As $A$ is a normal matrix, it can be diagonalized with a unitary matrix $U$, which is a set of orthonormal eigenvectors such that $U^{*}U=I$, and a diagonal matrix $\Lambda$, which has the eigenvalues of $A$ as entries, i.e. $\Lambda=\mathrm{diag}(\lambda_{1},\dots,\lambda_{n})$ [58, pp. 101–103]. Therefore,
@@
However, Theorem 1 serves as a strong theoretical proxy for the stability behavior of the ROM. Since the projection $\Phi$ inherently removes the high-frequency modes associated with the largest eigenvalues of $C(\mathbf{u})$, the inequality $\rho(C_{r})<\rho(C)$ is expected to hold robustly in practice. The theorem thus provides the structural justification for the increased stability limits observed in numerical experiments, even in the presence of trajectory divergence.